# Benchmarking quantum computers

### v1

@@ Abstract
There are many different kinds of quantum computer benchmark, and a bewildering array of specific benchmarks. We think that they can be best categorized and understood by how they interact with integrated quantum computers. Contemporary (prototype) quantum computing systems consist of multiple layers of hardware and software that is often called the quantum computing “stack” (see Fig. 3) in analogy to “software stacks” in classical computing [30]. How a benchmark interacts with this stack defines what parts of the quantum computer it tests, what performance metrics it measures, and whether the benchmark measures those metrics effectively.
@@
Quantum computing systems can be divided into interacting subsystems in many ways, but in our context it is most useful to adopt a stack that reflects the flow of information though a computation (Fig. 3). Although different quantum computing systems can have very different architectures, most systems’ information flow is approximately as follows. A quantum computing system’s input is a computational problem (e.g., “factor 15”) and its output is classical data that is intended to solve this problem (e.g., $3\times 5=15$). The output is computed by running quantum programs on qubits, wrapped within some potentially complex classical computing systems (that may be automated or human-aided).
@@
A quantum computer’s classical computing systems select an algorithm and then realize it as a quantum program. Here, a quantum program means a set of quantum circuits (see Box 2) that are perhaps wrapped inside a classical program, as in hybrid algorithms [49, 50, 51]. That quantum program is typically a high-level quantum program, i.e., it is written in a high-level language and its quantum circuits contain operations that are not native to that system’s computational qubits (which could be physical qubits or logical qubits built out of physical qubits running QEC). For example, it might use circuits that contain two-qubit gates between any pair of qubits (many systems have restricted connectivity), or large subroutines like the $n$-qubit quantum Fourier transform. This high-level program therefore typically needs to be compiled into a low-level quantum program whose circuits contain only gates that are “native” to that system’s computational qubits, before it can be executed. Finally, the system executes that low-level program, and processes the data. This conceptualization of the stack is an idealization, but it is a powerful aid for understanding and designing benchmarks.
@@
Good benchmarks have a well-defined procedure and this requires that their interactions with the stack be precisely stated. Most existing benchmarks’ interactions with the stack fall within the following simple schema shown in Fig. 3. The benchmark inserts computational tasks into one layer of the stack (the input layer) and those tasks propagate down the stack until they reach the benchmark’s output layer. That output is then analyzed to compute performance metrics. For example, the quantum volume benchmark inserts high-level circuits (which the system compiles and runs) and it analyzes the resultant data to estimate the system’s quantum volume [36]. Some benchmarks also specify limitations on what each layer of the stack is allowed to do (benchmark rules in Fig. 3), and can even modify the intermediate outputs of the stack (output modifiers in Fig. 3). These esoteric aspects of a benchmark can be critical to its robustness or efficiency, e.g., output modifiers can circumvent the problem of efficiently verifying the outputs of general quantum circuits (see Box 3).
@@
A benchmark quantifies the integrated performance of all the computational stages in between (and including) its input and output layers. Benchmarks that enter and exit the stack at different points quantify fundamentally different aspects of performance, and they constitute different and complementary categories of benchmark. Any entry and exit point in the stack defines a category of benchmarks. For example, compiler benchmarks [86, 87] isolate and test compilation algorithms by inserting high-level quantum circuits into the stack, and then analyzing the low-level circuits the compiler produces (using metrics such as the number of gates in those circuits). Three other categories of benchmark are also shown in Fig. 3, and are discussed below.
@@
A quantum computing system’s problem solving capabilities can be directly quantified using computational problem benchmarks that (i) challenge it to solve computational problems and then (ii) quantify performance in terms of metrics for the quality of (and/or time to) the solution. As illustrated in Fig. 3, these benchmarks enter the stack at the top level, and exit it at the lowest level, i.e., they test an entire computing system and compute a holistic performance metric. One reason such benchmarks are appealing is that they can be applied to any computing system, enabling comparison between radically different kinds of quantum computers as well as direct comparisons with classical computers.
@@
Any computational problem can be used to define a computational problem benchmark, but it is particularly appealing to use problems that are believed to be intractable for existing classical computers yet efficiently solvable with quantum algorithms. An example is the task “factor RSA-2048” with the binary performance metric “were the correct factors found?”. Benchmarks like this will become increasingly important once quantum utility is first achieved, but they provide few insights into the performance of today’s quantum computers that appear to be far from solving important classically hard problems (e.g., we can confidently predict that all of today’s systems will fail to factor RSA-2048). Computational problem benchmarks based on classically intractable problems must also address the problem of how to quantify the quality of a purported solution to that problem. This result verification is often challenging (see Box 3).
@@
Most of today’s algorithm-based benchmarks are not computational problem benchmarks [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 6, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72], i.e., they do not challenge a system to simply solve a computational problem like “factor 15”. This is because today’s quantum computers cannot solve useful classically intractable problems, and any classically tractable problem can be solved by a quantum computer using primarily (or even only) its classical computing resources. This is illustrated by a series of high-profile experiments that factored 15 using a compilation of Shor’s algorithm that was only possible because 15’s factors were already known [88]. So, although quantum computing hardware capabilities are often demonstrated by running quantum algorithms on simplified problem instances [89, 90, 91], these experiments are rarely intended to define formal benchmarks. A good benchmark based on solving classically tractable problems must restrict the kinds of classical computation that a quantum computer is allowed to use when running the benchmark (see Box 4). For this reason, most algorithm-based benchmarks that are used today are high- or low-level program benchmarks, which are two categories of benchmark that we now discuss.
Box 3. The verification problem in benchmarking. Most benchmarks quantify how well a quantum computer executed its tasks, but designing benchmarks that measure such metrics efficiently is difficult in general. Computational problem benchmarks (see Fig. 3) need to quantify the accuracy or correctness of a purported solution to their problem(s). The correctness of the solutions to some classically intractable problems (those in the co-NP complexity class, like factoring) can be easily validated or falsified, but this is atypical of problems with known quantum speedups. In those cases, natural metrics for solution quality will not typically be efficient to measure, and so a good benchmark must instead measure an efficient-to-estimate and reliable proxy for those metrics. This was the approach taken in Google’s random circuit sampling (RCS) benchmark experiments [2]. Program and circuit benchmarks (see Fig. 3) face a similar verification challenge: many natural metrics for quantifying how well a quantum computer executed a quantum circuit compare the observed and error-free outcome distributions of that circuit [60, 84]. However, directly calculating these metrics requires classically computing the error-free outcome distribution, which is exponential expensive in the number of qubits in general. Many benchmarks avoid this problem by using circuits that are efficiently simulable classically [20, 22, 23, 24, 25]. However this approach is typically inappropriate for high-level program benchmarks, because efficiently simulable circuits are often particularly easy to compile into shallow (or even trivial) circuits (see Box 4). One solution to this challenge is to define high-level program benchmarks using circuits that are potentially intractable to simulate classically, and to then “intercept” the compiled low-level circuits before they are run (see output modifiers in Fig. 3), replacing them with proxies for those circuits whose performance is provably similar but that are efficiently simulable [84, 92, 93].
@@
Quantum computations are implemented by running quantum programs, and so many benchmarks directly test a systems’ program-running capabilities. They typically do so by (i) tasking a quantum computer with running a set of quantum programs, and (ii) computing a metric that quantifies the error in those program’s execution (Box 3 discusses the challenge of efficiently computing such metrics). Because the core of any quantum program is the execution of quantum circuits, many of these benchmarks’ tasks are simply quantum circuits rather than more complex programs. Program (or circuit) benchmarks vary in the amount of compilation they allow (see Box 4), i.e., how high up they enter the stack, and they can be broadly categorized as high-level or low-level program benchmarks depending on whether they specify high-level programs that are then compiled by the system being tested, or low-level programs that permit little or no compilation.
@@
High-level program benchmarks (see Fig. 3) permit broad optimizations (compiling, routing, etc) of their programs or circuits. The most widely-used such benchmark is the quantum volume benchmark, which quantifies a system’s performance on random circuits containing random two-qubit gates coupling arbitrary pairs of qubits [36]. In contrast, many of the recently-developed high-level program benchmarks are based on applications and algorithms and therefore use highly structured programs [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 6, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Prominent examples include the QED-C’s benchmarking suite [60, 52] and SuperMarQ [55]. These suites consist of an assortment of benchmarks each of which is based on an algorithm, including algorithms that are illustrative but unlikely to be useful (e.g., the Bernstein–Vazirani algorithm) [60], heuristic NISQ algorithm [55, 67], algorithms that are likely to require a fault-tolerant architecture to provide utility (e.g., Shor’s algorithm) [60], and algorithms that prepare canonical quantum states like GHZ states [55].
@@
High-level program benchmarks are increasingly being used to compare different quantum computers [55], because they appear to enable simple and fair comparisons of systems with very different architectures. However, these comparisons can easily mislead. Contemporary quantum computers have not achieved quantum utility, so assessing these systems’ program-executing capabilities is useful only in so much as it assesses or incentivizes progress towards the goal of quantum utility. But improved performance on high-level program benchmarks can be obtained via system improvements that are unlikely to bring quantum utility closer. For example, many algorithms are unlikely to provide utility without fault tolerance, so an improvement to a quantum computing system’s classical compilation algorithm that enables better NISQ implementations of those algorithms (and therefore improves performance on existing algorithmic benchmarks) does not indicate progress towards quantum utility. The development of high-level program benchmarks that can reliably quantify, incentivize, and compare the progress of disparate architectures (e.g., NISQ versus fault-tolerant architectures) towards computational utility would be extremely valuable for the field.
@@
Low-level program benchmarks (see Fig. 3) forbid intrusive classical compilation of their programs, and this prevents improvements on these benchmarks being driven by unimportant compilation algorithm optimizations. Examples of low-level program benchmarks include many RB methods [17, 18, 19, 20, 22, 23, 24, 25, 26, 27, 28, 83], XEB [85], mirror circuit benchmarks [24], and some algorithm-based benchmarks. Low-level program benchmarks typically permit at most localized compilation of their circuits, meaning replacing each individual gate in a circuit with sequences of native gates that synthesize that gate (this kind of limited compilation, discussed in Box 4, is often described in terms of compilation barriers between circuit layers). Low-level program benchmarks directly quantify the performance of a system’s qubits and gates, so they are they valuable for discovering or quantifying unaccounted-for errors in qubits and gates, by comparing actual performance with that predicted by a model [37, 75, 24]. Some low-level program benchmarks compute holistic performance metrics, such as capability regions [37] (see Fig. 4) that directly quantify circuit execution error, but many of them are designed to extract error rates for individual logic operations.
Box 4. The role of compilation in benchmarking. Quantum programs can be written at many levels of abstraction, and the level of abstraction used in a benchmark’s programs fundamentally impacts what it measures. The core of any quantum program is one or more quantum circuits, and a circuit is expressed using gates from some set (a.k.a. basis). This gate set could contain high-level gates (e.g., the $n$-qubit quantum Fourier transform), or only canonical one- and two-qubit gates (like CNOT and Hadamard gates), or only a specific system’s low-level gates (e.g., cross-resonance gates between pairs of connected qubits). Benchmarks often allow their circuits to be compiled into different gate sets, as this is necessary to make a benchmark executable on many different systems. A quantum program refers (often implicitly) to an equivalence class of programs, i.e., “run this program” means “run any program in this program’s family”, and a well-defined high- or low-level program benchmark specifies the equivalence class for each of its programs. Integrated quantum computers can contain powerful classical computers, and effective, robust benchmarks must take this into account when defining the kinds of compilation that are allowed. Many benchmarks are highly permissive, e.g., each of the quantum volume benchmark’s circuits $C$ can be replaced by any circuit $C\textquoteright$ ’ that (in the absence of errors) implements (approximately) the same unitary as $C$. This is a common approach, as benchmarks like this jointly test a system’s compilation algorithms and qubits. A similar and conceptually simpler approach is for a benchmark to allow its circuits to be replaced with any other circuits that (in the absence of error) produce samples from the same probability distribution, but such benchmarks can be gamed. This is because (i) the benchmark permits replacing each of its circuits $C$ with a potentially trivial circuit $C^{\prime}$ that simply encodes the already-computed output distribution of $C$, and (ii) this is feasible to do (i.e., a system’s classical compilers could find such a $C^{\prime}$) whenever each $C$ can be quickly simulated classically.
@@
Many of the most mature benchmarks are designed to measure the error rates of the fundamental logic operations from which quantum circuits are built (e.g., individual single-qubit and two-qubit gates, layers of gates, measurements, etc). These component benchmarks (see Fig. 2) typically achieve this by running low-level circuits containing the components of interest and inferring how those components performed from the performance of those circuits. The paradigmatic component benchmarks are the RB protocols [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
@@
RB runs low-level circuits containing random gates, which average out the details of those gate’s error processes. This causes the success rate of these circuits to decay exponentially in circuit depth, and fitting this data to an exponential enables estimating those gates’ mean error rate. The de facto standard RB method measures the mean error rate of the set of all one- or two-qubit Clifford gates [20], and there is now a large family of RB methods and closely-related techniques, which adapt or improve standard RB in various ways. Prominent examples include XEB [85], direct RB [21], mirror RB [24, 25], binary RB [22], interleaved RB [29], cycle benchmarking [83], and character RB [27, 28]. XEB, direct RB, mirror RB, and binary RB are closely-related methods for measuring the average error rate of a set of layers of native gates. Interleaved RB and cycle benchmarking are methods for estimating the error rate of a single gate or layer of gates. Character RB enables adapting standard RB to sets of gates that form groups other than the Clifford group.
@@
Benchmarks that can measure the performance of fundamental operations on logical qubits will soon become increasingly important. Many of today’s benchmarks can be applied to logical qubits [26], but most of them were primarily designed and analyzed for benchmarking physical qubits. The degree to which the theory and assumptions underpinning existing benchmarks (e.g., approximate Markovianity [39]) will apply to logical qubits is currently unknown. However, we anticipate that new component benchmarks that are better suited to the properties of logical qubits will be needed. For example, some operations that logical qubits will perform are fundamentally different to any operations performed on physical qubits (e.g., lattice surgery [94]) and benchmarks that measure the performance of these complex computational primitives will be needed.
@@
The move towards fault-tolerant architectures also brings immediate needs for component benchmarks that focus on the properties of physical qubits and gates that are of most relevance in this setting. Many benchmarks measure the total rate of errors in fundamental logic operations (e.g., layers of gates), which is predictive of the failure rates of circuits executed directly on physical qubits and is therefore particularly relevant in a NISQ setting. However, some kinds of errors are much more damaging than others in fault-tolerant architectures, because they are more costly to correct. Similarly, some fundamental logical operations and subroutines are central to fault-tolerant quantum computing but of little or no importance in a NISQ architecture. Important examples include mid-circuit measurements, parity checks, and syndrome extraction cycles. Benchmarks for these primitives are relatively underdeveloped. Recent breakthroughs demonstrating components of fault-tolerant quantum computing [3, 95, 96, 97] introduced some prototype benchmarks, but mature benchmarks for the physical primitives of fault-tolerant quantum computing are a near-term need for the field.
@@
We think the most compelling current purpose of quantum computer benchmarks is measuring progress toward quantum utility. Correctly designed and interpreted benchmarks can help stakeholders understand and quantify the value of steps toward that goal. However, “quantum utility” is not a uniquely defined goal. There are many important, as-yet-unsolved problems that a quantum computer might solve, and many possible roadmaps to creating such a computer. We expect tracking progress to require a range of complementary benchmarks that are motivated by challenge problems, and interpreted relative to resource estimates and hardware roadmaps.
A challenge problem is a specific instance of an important computational problem that, if solved by a quantum computer, would constitute quantum utility. Challenge problems turn “quantum utility” into concrete computational goals. A resource estimate is a precise accounting of the computational resources required to solve a challenge problem. Resource estimates map computational goals to engineering goals. A hardware roadmap comprises plans and schedules for a sequence of increasingly capable quantum computers culminating in one that can solve a challenge problem. Hardware roadmaps define paths along which progress is tracked.
These ingredients enable choosing, creating, or adapting specific benchmarks that can test whether each milestone on a roadmap has been achieved, measure progress between milestones, and (perhaps) forecast the cost of future progress. In their absence, benchmarking cannot effectively track progress. Benchmark results can show that some property has improved, or that one device exceeds another in some way, but they can’t quantify the value of that change.

### v2

@@ Abstract
There are many different kinds of quantum computer benchmark, and a bewildering array of specific benchmarks. We think that they can be best categorized and understood by how they interact with integrated quantum computers. Contemporary (prototype) quantum computing systems consist of multiple layers of hardware and software that is often called the quantum computing “stack” (see Fig. 3) in analogy to “software stacks” in classical computing [34]. How a benchmark interacts with this stack defines what parts of the quantum computer it tests, what performance metrics it measures, and whether the benchmark measures those metrics effectively. In this section, we explain this framework for understanding benchmarks. We then discuss three kinds of benchmark that are, in our view, of particularly broad importance.
@@
Quantum computing systems can be divided into interacting subsystems in many ways, but in our context it is most useful to adopt a division that reflects the flow of information though a computation (Fig. 3). Although different quantum computing systems can have very different architectures, most systems’ information flow is approximately as follows. A quantum computing system’s input is a computational problem (e.g., “factor 15”) and its output is classical data that is intended to solve this problem (e.g., $3\times 5=15$). The output is computed by running quantum programs on qubits, wrapped within some potentially complex classical computing systems (that may be automated or human-aided).
@@
A quantum computer’s classical computing systems select an algorithm and then realize it as a quantum program. Here, a quantum program means a set of quantum circuits (see Box 2) that are perhaps wrapped inside a classical program, as in hybrid algorithms [63, 64, 65]. That quantum program is typically a high-level quantum program, i.e., it is written in a high-level language and its quantum circuits contain operations that are not native to that system’s computational qubits (which could be physical qubits or logical qubits built out of physical qubits running QEC). For example, it might use circuits that contain two-qubit gates between any pair of qubits (many systems have restricted connectivity), or large subroutines like the $n$-qubit quantum Fourier transform. This high-level program therefore typically needs to be compiled into a low-level quantum program whose circuits contain only gates that are “native” to that system’s computational qubits, before it can be executed. Finally, the system executes that low-level program, processes the data (which can include complex processing like error mitigation [7, 101]) and returns results. This conceptualization of the stack is an idealization, but it is a powerful aid for understanding and designing benchmarks.
@@
A benchmark should have a well-defined procedure, and this requires that its interactions with the stack are precisely stated. Most existing benchmarks’ interactions with the stack fall within the following simple schema shown in Fig. 3. The benchmark inserts computational tasks into one layer of the stack (the input layer) and those tasks propagate down the stack until they reach the benchmark’s output layer. That output is then analyzed to compute performance metrics. For example, the quantum volume benchmark inserts high-level circuits (which the system compiles and runs) and it analyzes the resultant data to estimate the system’s quantum volume [52]. Some benchmarks also specify limitations on what each layer of the stack is allowed to do (benchmark rules in Fig. 3), and can even modify the intermediate outputs of the stack (output modifiers in Fig. 3). These esoteric aspects of a benchmark can be critical to its robustness or efficiency, e.g., output modifiers can circumvent the problem of efficiently verifying the outputs of general quantum circuits (see Box 4).
@@
A benchmark quantifies the integrated performance of all the computational stages in between (and including) its input and output layers. Benchmarks that enter and exit the stack at different points quantify fundamentally different aspects of performance, and they constitute different and complementary categories of benchmark. Any entry and exit point in the stack defines a category of benchmarks. Some categories of benchmarks test only classical components of a quantum computer, such as compilers [102, 103] or qubit control electronics [104]. For example, compiler benchmarks [102, 103] isolate and test compilation algorithms by inserting high-level quantum circuits into the stack, and then analyzing the low-level circuits the compiler produces (using metrics such as the number of gates in those circuits). However, in our view, the benchmarks of most broad interest and importance test the computational qubits and gates. Three such categories of benchmark are shown in Fig. 3, and we discuss these kinds of benchmark in the remainder of this section.
@@
A quantum computing system’s problem solving capabilities can be directly quantified using computational problem benchmarks that challenge it to solve computational problems and quantify performance in terms of metrics for the quality of (and/or time to) the solution. As illustrated in Fig. 3, these benchmarks enter the stack at the highest level, and exit it at the lowest level, i.e., they jointly test an entire computing system. One reason these benchmarks are appealing is that they can be applied to any computing system, enabling comparison between radically different kinds of quantum computers, as well as direct comparisons with classical computers.
@@
Any computational problem can be used to define a computational problem benchmark, but it is appealing to use problems that are believed to be intractable for classical computers yet efficiently solvable with quantum algorithms. An example is the task “factor RSA-2048” with the binary performance metric “were the correct factors found?” Benchmarks like this will become increasingly important once quantum utility is first achieved, although note that such benchmarks must address the often-challenging problem of how to quantify the quality of a purported solution to a classically intractable problem (see Box 4).
@@
Computational problem benchmarks that use practical, classically intractable problems provide few insights into the performance of today’s quantum computers, because they appear to be far from solving such problems (see Fig 3). This has motivated benchmarking quantum computers’ abilities to solve classically tractable problems (e.g., “factor 15”), but benchmarks of this sort can easily mislead. This is because any such problem can be solved by a quantum computer using primarily (or even only) its classical computing resources. This is illustrated by a series of high-profile experiments that factored 15 using a compilation of Shor’s algorithm that was only possible because 15’s factors were already known [105]. Quantum computing hardware capabilities are often demonstrated by running quantum algorithms on simplified problem instances [106, 107, 108], but these experiments are rarely intended to define formal benchmarks. A good benchmark based on solving classically tractable problems must restrict the kinds of classical computation that a quantum computer is allowed to use when running the benchmark (see Box 3). For this reason, most algorithm-based benchmarks that are used today [66, 53, 67, 68, 69, 70, 71, 72, 73, 74, 75, 7, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85] are best categorized as high- or low-level program benchmarks, rather than computational problem benchmarks, which are two categories of benchmark that we now discuss.
Box 3. The role of compilation in benchmarking. Quantum programs can be written at many levels of abstraction, and the level used in a benchmark’s programs fundamentally impacts what it measures. The core of any quantum program is one or more quantum circuits, and a circuit is expressed using gates from some set, also known as a basis. This gate set could contain high-level gates (e.g., the $n$-qubit quantum Fourier transform), or only canonical one- and two-qubit gates (e.g., the CNOT and Hadamard gates), or only a particular system’s low-level gates (e.g., cross-resonance gates between pairs of connected qubits). However, a benchmark will only be executable on a broad range of quantum computing systems if it permits its circuits to be compiled into circuits over different gate sets (we use “compilation” to mean the process of turning one circuit into another circuit that implements the same overall computation, which is sometimes subdivided into various interrelated tasks such as gate synthesis, transpilation, and routing). More generally, a quantum program refers to an equivalence class of programs, i.e., “run this program” means “run any program in this program’s family”, and a well-defined benchmark whose tasks are quantum programs specifies the equivalence class for each of its programs. Integrated quantum computers can contain powerful classical computers, and effective, robust benchmarks must take this into account when defining the kinds of compilation that are allowed. This is illustrated by the de facto standard RB method [24]. $n$-qubit standard RB measures the mean error rate of a set of high-level $n$-qubit quantum circuits, the set of all $n$-qubit Clifford unitaries, that can be compiled in any way. Standard RB measures the “error-per-Clifford” by running circuits consisting of $d+1$ random Clifford unitaries (with $d$ varied) that, if composed together, implements the trivial, identity operation. This enables precise estimation of the error-per-Clifford, but only if sequential Clifford unitaries are not compiled together—which is typically described using compilation barriers between each of the Clifford unitaries. In our framework of Fig. 3, standard RB is a high-level program benchmark that (like many useful benchmarks) specifies rules that limit what the compiler is allowed to do. Many high-level program benchmarks are highly permissive in the kinds of compilations they allow, e.g., each of the quantum volume benchmark’s circuits $C$ can be replaced by any circuit $C$’ that, in the absence of errors, implements approximately the same unitary as $C$. This is a common approach, as benchmarks like this jointly test a system’s compilation algorithms and qubits. A similar and conceptually simpler approach is for a benchmark to allow its circuits to be replaced with any other circuits that (in the absence of error) produce samples from the same probability distribution, but such benchmarks can be gamed. This is because the benchmark permits replacing each of its circuits $C$ with a potentially trivial circuit $C^{\prime}$ that simply encodes the already-computed output distribution of $C$, and this is feasible to do (i.e., a system’s classical compilers could find such a $C^{\prime}$) whenever each $C$ can be quickly simulated classically.
@@
Quantum computations are implemented by running quantum programs, and so many benchmarks directly test a system’s program-running capabilities. They typically do so by (i) tasking a quantum computer to run a set of quantum programs, and (ii) computing a metric that quantifies the error in those programs’ execution (Box 4 discusses the challenge of efficiently computing such metrics). Because the core of any quantum program is the execution of quantum circuits, many of these benchmarks’ tasks are simply quantum circuits rather than more complex programs. Program (or circuit) benchmarks vary in the amount of compilation they allow (see Box 3), i.e., how high up they enter the stack, and they can be broadly categorized as high-level or low-level program benchmarks.
Program benchmarks also vary in where they exit the stack. Some program benchmarks (such as RB) specify analysis of the raw bit string data from circuits. Some explicitly permit data post-processing including complex error mitigation [66]. Others are ambiguous about what, if any, post-processing is allowed. Our view is that the results of benchmarks run with error mitigation should currently be interpreted with extreme caution, because clear connections between the error-mitigated performance of today’s prototype quantum computers and progress towards quantum utility have not been made.
@@
High-level program benchmarks (see Fig. 3) permit broad optimizations (compiling, routing, etc) of their programs or circuits. The most widely-used such benchmark is the quantum volume benchmark, which quantifies a system’s performance on random circuits containing random two-qubit gates coupling arbitrary pairs of qubits [52]. In contrast, many of the recently-developed high-level program benchmarks are based on applications and algorithms and therefore use highly structured programs [66, 53, 67, 68, 69, 70, 71, 72, 73, 74, 75, 7, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. Prominent examples include the QED-C’s benchmarking suite [73, 66] and SuperMarQ [68]. These suites consist of an assortment of benchmarks each of which is based on an algorithm, including algorithms that are illustrative but unlikely to be useful (e.g., the Bernstein–Vazirani algorithm) [73], heuristic NISQ algorithms [68, 80], algorithms that are likely to require a fault-tolerant architecture to provide utility (e.g., Shor’s algorithm) [73], and algorithms that prepare canonical quantum states like GHZ states [68].
@@
High-level program benchmarks are increasingly being used to compare different quantum computers [6, 73, 68, 81, 80, 70], and it has been claimed that this enable simple and fair comparisons of systems with very different architectures. In our view, these comparisons can easily mislead. Contemporary quantum computers have not achieved quantum utility, so our perspective is that assessing these systems’ program-executing capabilities is useful only in so much as it assesses or incentivizes progress towards the goal of quantum utility. But improved performance on high-level program benchmarks can be obtained via system improvements that are unlikely to bring quantum utility closer. For example, many algorithms are unlikely to provide utility without fault tolerance, so an improvement to a quantum computing system’s classical compilation algorithm that enables better NISQ implementations of those algorithms (and therefore improves performance on existing algorithmic benchmarks) does not indicate progress towards quantum utility. The development of high-level program benchmarks that can reliably quantify, incentivize, and compare the progress of disparate architectures (e.g., NISQ versus fault-tolerant architectures) towards computational utility would be extremely valuable for the field.
@@
Low-level program benchmarks (see Fig. 3) forbid intrusive classical compilation of their programs. This prevents improvements on these benchmarks being driven by compilation algorithm optimizations. Examples of low-level program benchmarks include many RB methods [21, 22, 23, 24, 26, 27, 28, 29, 30, 31, 32, 56], XEB [55], mirror circuit benchmarks [28], and some algorithm-based benchmarks. Low-level program benchmarks typically permit only localized compilation of their circuits, i.e., each individual gate in a circuit can be replaced with a sequence of native operations that synthesize that gate (see Box 3). Low-level program benchmarks directly quantify the performance of a system’s qubits and gates, so they enable discovery and quantification of unexpected errors in qubits and gates, by comparing actual performance to that predicted by a model [39, 91, 28]. Some low-level program benchmarks compute holistic performance metrics, such as capability regions [39] (see Fig. 4) that directly quantify circuit execution error, but many of them are designed to extract error rates for individual logic operations.
Box 4. The verification problem in benchmarking. Most benchmarks quantify how well a quantum computer executed its tasks, but designing benchmarks that measure such metrics efficiently is difficult in general. Computational problem benchmarks (see Fig. 3) need to quantify the accuracy or correctness of a purported solution to their problem(s). The correctness of the solutions to some classically intractable problems (those in the co-NP complexity class, like factoring) can be easily validated or falsified, but this is atypical of problems with known quantum speedups. In those cases, natural metrics for solution quality will not typically be efficient to measure [109, 110], and so a good benchmark must instead measure an efficient-to-estimate and reliable proxy for those metrics. Google’s famous random circuit sampling (RCS) experiments utilized a procedure of this sort, specialized to the RCS problem [3]. More general methods for verifying the outputs of quantum computations have been developed [109, 110, 37, 35, 111, 112], including approaches based on cryptographic techniques [111] or comparisons to a trusted quantum computer [113]. However, some of these methods make strong assumptions about the nature of a quantum computer’s errors. Program and circuit benchmarks (see Fig. 3) face a similar verification challenge: many natural metrics for quantifying how well a quantum computer executed a quantum circuit compare the observed and error-free outcome distributions of that circuit [73, 100]. However, directly calculating these metrics requires classically computing the error-free outcome distribution, which is exponential expensive in the number of qubits in general. Many benchmarks avoid this problem by using circuits that are efficiently simulable classically [24, 26, 27, 28, 29]. However this approach is typically inappropriate for high-level program benchmarks, because efficiently simulable circuits are often particularly easy to compile into shallow (or even trivial) circuits (see Box 3). One solution to this challenge is to define high-level program benchmarks using circuits that are potentially intractable to simulate classically, and to then “intercept” the compiled low-level circuits before they are run (see output modifiers in Fig. 3), replacing them with proxies for those circuits whose performance is provably similar but that are efficiently simulable [100, 114, 115].
@@
Many of the most mature benchmarks are designed to measure the error rates of the fundamental logic operations from which quantum circuits are built (e.g., individual single-qubit and two-qubit gates, layers of gates, measurements, etc). These component benchmarks (see Fig. 2) typically achieve this by running low-level circuits containing the components of interest and inferring how those components performed from the performance of those circuits. RB protocols [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] are the paradigmatic component benchmarks.
@@
RB methods run low-level circuits containing random gates, which average out the details of those gates’ errors. This causes the mean success rate of these circuits to decay exponentially in circuit depth, at a rate equal to the mean of those gates’ average-case error rate (i.e., process infidelity, or, equivalently, average gate infidelity [1]). The de facto standard RB method measures the error rate of the set of all one- or two-qubit Clifford gates [24]. There are now many RB methods and closely-related techniques that adapt or improve standard RB. Prominent examples include direct [25], mirror [28, 29], binary [26], character [31, 32], filtered [32, 116], and interleaved RB [33], as well as XEB [55] and cycle benchmarking [56]. XEB, direct RB, mirror RB, and binary RB are closely related methods for measuring the average error rate of a set of native gate layers. Interleaved RB and cycle benchmarking are methods for estimating the error rate of a single gate or layer of gates. Character RB and other filtered RB methods extend RB to sets of gates that are (or generate) groups other than the Clifford group (or another unitary 2-design).
Tomographic characterization protocols [45, 46, 47, 48, 49, 1] are sometimes repurposed to benchmark components. The primary purpose of tomography protocols is to estimated detailed, predictive error models. For example, quantum process tomography uses a lot of data to estimate the $4^{n}\times 4^{n}$ process matrix describing a gate or subroutine [45, 49, 1]. A tomographically-estimated model does not constitute a performance metric, but summary performance metrics such as gate fidelity can be extracted from them, and this constitutes a benchmark. This approach is very data-intensive, but once tomography is complete, essentially any metric can be estimated. This includes metrics that no known benchmark can measure efficiently, including worst-case error rates such as diamond norm error [36]. Worst-case metrics are valuable, but are rarely reported because they are pessimistic and grow rapidly harder to measure with the number of qubits $n$. Tomography-derived benchmarks are rarely used for more than two qubits.
@@
Benchmarks that can measure the performance of fundamental operations on logical qubits will soon become increasingly important. Most of today’s benchmarks can be applied to logical qubits [30] but they were primarily designed for physical qubits. Whether the theories and assumptions underpinning existing benchmarks (e.g., approximate Markovianity [40]) will apply to logical qubits is largely unknown. However, we anticipate that new component benchmarks that are designed for logical qubits will be needed, e.g., because some logical qubit operations (such as lattice surgery [117]) are fundamentally different from any operations on physical qubits.
@@
The move towards fault-tolerant architectures also brings immediate needs for component benchmarks that measure the properties of physical qubits and gates that are most relevant in this setting. Many benchmarks measure the total rate of errors in layers of gates [25, 28, 29, 26, 27, 56], which can predict the failure rates of circuits executed directly on physical qubits (e.g., NISQ algorithms). However, in fault-tolerant architectures some kinds of errors are much more damaging than others, because they are more costly to correct [15]. Similarly, some physical operations and subroutines are central to fault-tolerant quantum computing but are not essential in NISQ architectures. Important examples include mid-circuit measurements, parity checks, and syndrome extraction cycles [4, 18, 19, 20]. Some prototype benchmarks exist for the physical primitives of fault-tolerant quantum computing [118, 119, 120, 4, 18, 19, 20], but mature benchmarks are a near-term need for the field.
@@
We conclude our perspective by discussing what is, in our view, the most compelling current and near-term purpose of quantum computer benchmarks: measuring progress toward quantum utility. The single biggest challenging here is that “quantum utility” is not a uniquely defined goal. There are many important problems that a quantum computer might solve, and many possible routes to creating such a computer. Our view is that tracking progress towards quantum utility will require a range of complementary benchmarks that are motivated by challenge problems (computational problems that, if solved by a quantum computer, would constitute quantum utility), and interpreted relative to resource estimates (an accounting of the computational resources required to solve a challenge problem) and hardware roadmaps (plans for how to build a quantum computer that can solve a challenge problem). In this section, we discuss challenge problems, resource estimates, and hardware roadmaps—summarizing each area from the perspective of how it impacts benchmarking—and then we conclude by presenting our views on how benchmarks can measure progress towards quantum utility.